OpenStudy (anonymous):

whats the reference angle of 13pie/9.whats the supplementary angle to the reference angle of this.

OpenStudy (anonymous):

the reference angle is 260. the supplementary angle is 100. i'm not so sure..

OpenStudy (anonymous):

i have the degree angle is 260

OpenStudy (anonymous):

degree measure i emean

OpenStudy (anonymous):

the reference angle is 13pi/9 - pi or in short \[\frac{13\pi}9 - \frac{9\pi}9\] the supplementary angle is just pi minus that

OpenStudy (anonymous):

how did u get the -9pie/9